### v1

# Learning Semantic Traversability with Egocentric Video and Automated Annotation Strategy

Abstract
For reliable autonomous robot navigation in urban settings, the robot must have the ability to identify semantically traversable terrains in the image based on the semantic understanding of the scene. This reasoning ability is based on semantic traversability, which is frequently achieved using semantic segmentation models fine-tuned on the testing domain. This fine-tuning process often involves manual data collection with the target robot and annotation by human labelers which is prohibitively expensive and unscalable. In this work, we present an effective methodology for training a semantic traversability estimator using egocentric videos and an automated annotation process. Egocentric videos are collected from a camera mounted on a pedestrian’s chest. The dataset for training the semantic traversability estimator is then automatically generated by extracting semantically traversable regions in each video frame using a recent foundation model in image segmentation and its prompting technique. Extensive experiments with videos taken across several countries and cities, covering diverse urban scenarios, demonstrate the high scalability and generalizability of the proposed annotation method. Furthermore, performance analysis and real-world deployment for autonomous robot navigation showcase that the trained semantic traversability estimator is highly accurate, able to handle diverse camera viewpoints, computationally light, and real-world applicable. The summary video is available at https://youtu.be/EUVoH-wA-lA.
Index Terms:
Vision-based Navigation, Semantic Scene Understanding, Deep Learning for Visual Perception
I Introduction
Mobile robots should possess autonomous navigation capabilities as it is an essential feature for various applications, including package delivery, factory inspection, urban security, and search and rescue. In the realm of autonomous navigation, various problem domains exist. Unlike the majority of prior research, which primarily addresses autonomous off-road navigation in field environments[1, 2, 3, 4, 5] or autonomous on-road navigation in urban settings[6, 7], our focus lies in tackling the challenges of autonomous off-road navigation within urban environments.
When people navigate in an urban environment, they usually start by utilizing readily available map services on their smartphones, such as “Google Maps”, to find a rough global route from their current location to their desired destination. They then repeat the process of planning and tracking a local path within their field of view that is most similar to the global path. To ensure safe and responsive closed-loop navigation, not only the similarity with the global path is taken into account when determining the local path, but also the geometric and semantic traversability. Geometric traversability is evaluated using geometric information from the current viewpoint to avoid collisions and rough, bumpy terrains. Semantic traversability is determined by semantic understanding and prior knowledge of the scene, which results in a preference for sidewalks over roads and the use of crosswalks when crossing the street. Humans’ navigation procedures and reasoning capabilities serve as valuable insights for developing an autonomous navigation framework for mobile robots.
Figure 1: Overall framework
Robot navigation frameworks are often built hierarchically with multiple modules, including a mapper, path planner, and controller. The mapper creates a map of the environment in a suitable representation using exteroceptive sensors such as lidar and camera. The path planner then finds a path on the map that maximizes the objective function defined by the engineer. The controller then follows the planned path by executing appropriate high-level (e.g., velocity commands, waypoints)[8] and low-level commands (e.g., joint PD targets)[9]. Similar to how humans navigate, a global route connecting the current location to the desired destination can be obtained from an openly accessible world map. Subsequently, the local path planner adjusts the computed path based on geometric and semantic traversability, which are evaluated and accumulated in the built map[10] during navigation. Geometric traversability is computed using either analytic techniques[11] or learning-based methods[8, 12] from a map constructed with depth cameras or lidar, which contains geometric information such as grid occupancy and terrain height. Semantic traversability is often evaluated using semantic segmentation predictions from semantic segmentation models[13, 14] trained on large-scale datasets containing RGB images and semantic class annotations. Engineers assign a heuristic objective value to each terrain class based on their preference[1, 15, 16]. For example, if the engineer wants the robot to have a preference $Sidewalk>Grass>Road$, the objective values for each terrain class can be set as $Sidewalk=1,\,Grass=0.5,\,Road=0.25$, indicating high values for high preference.
However, directly computing semantic traversability based on semantic segmentation predictions limits its performance to that of the trained semantic segmentation model. Although the deployed semantic segmentation models are trained on large-scale datasets, they often show limited generalization performance due to a significant domain gap between the training and testing datasets[17]. Specifically, the models are vulnerable to out-of-distribution (OOD) RGB images (e.g., a model trained on autonomous-driving datasets results in a strong bias toward the camera viewpoint of street vehicles and shows limited performance in non-vehicle regions such as pedestrian areas) and OOD terrain classes (e.g., a model trained to recognize $Grass,\,Sand,\,and\,\,Rock$ cannot be directly used to predict classes such as $Sidewalk,\,Road,\,and\,\,Crosswalk$).
Because incorrect classifications of terrain classes can deteriorate navigation behavior, the trained semantic segmentation models are usually fine-tuned before being deployed with manually collected and annotated data from the testing domain[16]. Concretely, the data is collected by rolling out the robot in the testing domain, often controlled by human experts, and the data annotation is held manually by humans. However, the process of manually preparing the semantic segmentation dataset is highly expensive, labor-intensive, and not scalable. For rapid development of the semantic traversability estimator, a more scalable data collection method and an automated labeling technique are necessary. Furthermore, the model design of the semantic traversability estimator should be adaptable enough to handle OOD terrain classes and OOD visual features.
To this end, we present an effective methodology for training a neural network that predicts semantic traversability using a highly scalable data acquisition method and an automated annotation strategy without the need for manual labeling (Figure 1). Instead of using the target robot to gather data, we utilize egocentric videos obtained by mounting a camera on a person’s chest and capturing the egocentric perspective of a pedestrian crossing an urban area. The collected egocentric videos are then automatically labeled using the Segment Anything Model (SAM)[18], a recent foundation model in image segmentation possessing high generalizability and powerful prompting techniques. The semantically traversable regions in the RGB image are annotated by prompting SAM with approximate footsteps extracted from the egocentric videos using the Monocular Visual-Inertial SLAM[19], eliminating the need for human intervention during the annotation and prompting process. The semantic traversability estimator is then obtained by fine-tuning a lightweight semantic segmentation model[13] with the auto-labeled dataset and minimal model architecture modification. We have conducted extensive quantitative and qualitative evaluations to demonstrate the effectiveness of our method in terms of data collection, data annotation, and model training. Specifically, we show that leveraging egocentric video with the proposed automated annotation strategy possesses great potential for highly scalable and worldwide training data collection due to its minimal hardware requirements. Additionally, we demonstrate that fine-tuning the pre-trained semantic segmentation model yields a semantic traversability estimator that is highly accurate, able to handle diverse camera viewpoints, and lightweight (71Hz on a desktop GPU and 16Hz on an onboard embedded GPU). The trained semantic traversability estimator is further deployed for autonomous quadruped robot navigation in an urban environment, showcasing the method’s capability for real-world applications.
II Related work
There has been various prior research on learning a geometric traversability estimator (typically parameterized with neural networks) in a self-supervised manner without the need for manual labels. The required raw data, such as RGB-D images, was first collected by rolling out the robot in the real-world testing area or the physics simulation[8, 12]. Then, traversability labels were obtained from the human expert’s velocity command[20], velocity command tracking performance[21], collision checking results[8, 12], or IMU sensor data[3, 5]. However, the methods are limited to geometric traversability and place less emphasis on category-level semantic information. Thus, the trained estimators cannot distinguish between terrains that are geometrically similar but semantically different, such as a sidewalk and a flat road.
Several methods have previously been proposed for training a semantic traversability estimator via self-supervised learning. Wellhausen et al.[2] used the future robot footsteps projected on the current image as traversable point labels for training the traversability estimator. Each point label included ground reaction scores from force-torque sensors as well as corresponding terrain classes. Schmid et al.[22] approached the problem as an anomaly detection problem, training an autoencoder to reconstruct only the traversed area of the image. The traversable region on the image was then identified as an area with a low reconstruction error. Frey and Mattamala et al.[4] integrated weak image segmentation masks with automatically derived traversability values from velocity command tracking errors, allowing the traversability estimator to consider semantic cues. There were methods focused on representation learning[23, 24] in which terrain feature encoders were first trained using unsupervised learning and then a binary terrain classifier was trained to determine whether the terrain is preferred or not based on the pre-trained feature. However, these methods necessitate a two-step training process and rely on manual data labeling or heuristics to generate binary terrain preference labels.
Our work is closely related to the ones presented by Wellhausen et al.[2] and Frey and Mattamala et al.[4] but differs from theirs in three ways. First, we obtained traversable area labels from each image by using the recent foundation model in image segmentation (i.e., SAM) and the model’s powerful prompting capability. The proposed annotation method results in large area labels, which are based on visual cues like colors and edges, instead of local point or segment labels. There is a concurrent work done by Jung et al.[25] that used SAM with contrastive learning in self-supervised traversability learning for navigation in field environments. Our work distinguishes itself by its emphasis on urban settings and the integration of semantic information in the annotation process. Second, we utilized egocentric videos taken with a camera mounted on a pedestrian’s chest instead of calibrated sensor modules mounted on the target robot. This minimal hardware requirement enables effective and scalable data collection. Lastly, we tested our approach in an urban setting, which presents a greater challenge for reliable traversability estimation compared to park-like settings or field environments because of the high visual diversity and significance of semantic information.
III Method
In this section, we first describe the data collection and automatic annotation pipeline for creating a dataset to train the semantic traversability estimator, followed by an elaboration on its training with the created dataset. We then illustrate how we integrate the trained semantic traversability estimator into the hierarchical navigation framework and use it for real-world robot autonomy in urban settings.
III-A Data collection
We collect egocentric videos obtained by mounting a GoPro camera on a person’s (i.e., volunteer) chest and capturing the egocentric perspective of the volunteer walking around an urban environment. This approach enables effective and scalable data collection compared to previous works[2, 4, 22, 24, 23] that rely on data obtained directly from the target robot. Leveraging the target robot itself often requires well-calibrated sensor settings and limits the data acquisition capability based on several constraints such as the number of available robots, the robot’s speed, and physically reachable area. The volunteer is instructed to walk around inside the designated area while taking into account their visual path preference, which is consistent with urban regulations such as pedestrians walking on the sidewalk rather than the road and using crosswalks when crossing the road. The volunteer is not restricted from going to specific terrains, such as steep bumps, because our target robot for deploying the semantic traversability estimator is a legged robot that can traverse most regions that humans can traverse. However, volunteers may be given instructions in advance to avoid certain terrains, taking into account the type of target robot and its controller capability, which prevents the corresponding terrain from being labeled as traversable. For example, if the target robot is a wheeled robot, the volunteer may be told to avoid stairs and vertical slopes because the robot cannot overcome them.
The methodology of leveraging egocentric videos obtained from humans for robot navigation is inspired by the Ego4D dataset[26], a massive-scale egocentric video dataset capturing diverse daily-life activities. The work[26] demonstrated the possibility of collecting egocentric videos on a large scale, and the following works showed the usage of egocentric videos in robotic tasks such as affordance learning for robot manipulation[27]. Our work aims to take advantage of egocentric videos in robot navigation and use them to learn semantic traversability.
III-B Automatic annotation
We develop an automatic annotation pipeline that extracts visually preferred regions to walk in each frame of the collected egocentric videos. First, ORB-SLAM3[19] is used to estimate the camera trajectory—a set of coordinates representing the camera’s position and orientation with respect to a fixed global reference frame—from each video clip. We can obtain global camera poses on an absolute scale by running the Monocular Visual-Inertial SLAM with the synchronized IMU sensor built inside the GoPro camera. Because there is not much of a visual difference between every two consecutive frames (in our case, the camera frame rate was 24Hz), we only use camera frames that are identified as keyframes. Let’s denote the estimated SE(3) pose trajectory of the camera in a single video clip as ${}^{W}\textbf{T}_{0:N}=[{}^{W}\textbf{R}_{0:N}|{}^{W}\textbf{P}_{0:N}]$, where ${}^{W}\textbf{R}$ and ${}^{W}\textbf{P}$ are each rotation matrix and position vector, $N$ is the number of detected keyframes, and $W$ is the fixed global reference frame that corresponds to the gravity-aligned frame at the initial camera pose. We then obtain the approximate terrain points that the volunteer traversed, utilizing the fact that the camera is mounted on the volunteer’s chest as
${}^{W}\textbf{P}_{0:N}^{f}={}^{W}\textbf{P}_{0:N}-[0,0,H],$ (1)
where $H$ is the measured distance from the ground to the mounted camera when the volunteer stands still. We will refer to these approximate terrain points that the volunteer traversed as footsteps. The footsteps are projected into the image space as
$\textbf{p}_{i:i+n_{i}}^{f}=\textbf{K}\cdot\textbf{T}_{CW}^{i}\cdot{}^{W}\textbf{P}_{i:i+n_{i}}^{f},$ (2)
where $C$ is the camera frame, K is the intrinsic camera calibration matrix, $\textbf{T}_{CW}^{i}={}^{W}\textbf{T}_{i}^{-1}$ is the extrinsic transformation matrix from the fixed global reference frame into the $i$th camera frame, and $n_{i}$ is the number of considered keyframes at the $i$th camera frame. The number of future keyframes $n_{i}$ to be taken into account for each frame is determined by the predefined time horizon $T^{\prime}$. In particular, keyframes between $t$ and $t+T^{\prime}$ time stamps are taken into consideration for the $i$th camera frame that corresponds to the $t$ time stamp. We ignore potential occlusions of footsteps caused by vertical geometry and dynamic obstacles, and we also note that a short time horizon $T^{\prime}$ (in our case, $T^{\prime}=3s$) resulted in almost no occlusion. In the future, occluded footsteps can be removed using a 3D reconstruction of the environment.
The semantically traversable area in the current frame that the volunteer intends to traverse is obtained from the projected footsteps. Our key insight is that the semantically preferred region in the image will include the projected footsteps and differ from the semantically undesirable region by visual cues like edges, colors, and textures. These visual cues are important for humans to distinguish between geometrically similar but semantically distinct paths (e.g., a road and a sidewalk, a bicycle track and a pedestrian track), and urban planning regulations also take them into account. To this end, we leverage SAM with its powerful prompting technique to extract the semantically desirable area. SAM, trained on an extremely large-scale dataset of 1 billion images, demonstrates high generalizability across diverse images and robust performance in segmenting regions that align well with given prompts. For each frame, we obtain mask predictions from SAM by passing the model with the corresponding image and projected footsteps as positive point prompts. Although SAM can handle both RGB and gray-scale images, we used gray-scale images since SAM’s results were highly color-sensitive. When projecting points from 3D space to image space, their distribution on the image becomes non-uniform, concentrating more densely in regions far from the observer. These non-uniform point prompts may degrade the predicted mask from SAM, and thus $n_{p}$ (in our case, $n_{p}=3$) points are sampled from the projected footsteps $\textbf{p}_{i:i+n_{i}}^{f}$ via farthest point sampling and used for positive point prompts. Additionally, we post-process the raw mask predictions from SAM with an area and contour filter. The area filter selects a mask from the proposals if its area exceeds a certain threshold. The contour filter extracts contours from mask predictions and leaves only the contour with the biggest area. This allows for the removal of tiny, independently segmented pieces.
Figure 2: Automatic annotation process
We use a large semantic segmentation model [14] in conjunction with SAM to produce the final semantic traversability labels. This process introduces more detailed semantic information required for urban environment deployment (e.g., road is less preferred compared to other traversable terrains) and can be removed if not necessary (e.g., park-like environment). Even though a large semantic segmentation model’s computational load makes it unsuitable for high-frequency real-time usage (i.e., Mask2former[14] model that we used runs at 1.66Hz on the onboard embedded GPU), this is less important for offline data annotation. From the semantic segmentation predictions, segments labeled as crosswalk and road are added to and removed from the SAM mask predictions, respectively. If projected footsteps are sufficiently included in the remaining masks, the remaining segmented areas are annotated with a value of 1. If not, this is typically a scenario when the volunteer walked on the road because it was the best navigation behavior in their current field of view. In this instance, the areas where the SAM mask intersects with areas identified as road are given a value of 0.25, whereas the remaining SAM mask areas and areas identified as crosswalk are given a value of 1. The annotation process is summarized in Figure 2.
III-C Model training
We use the dataset acquired through our automatic annotation pipeline to train a semantic traversability estimator. The dataset is composed of tuples of RGB images and pixel-wise traversability values ranging from 0 to 1, with higher values indicating a stronger preference. We leverage neural networks to parameterize the semantic traversability estimator. However, rather than training it from scratch, we model it with a lightweight semantic segmentation model trained on a large-scale dataset (i.e., Segformer[13]) and fine-tune it with the collected dataset. A single 1x1 convolution and sigmoid activation layer is added to Segformer’s final layer to reduce the number of channels from the number of segmentation classes to one and restrict the output from 0 to 1. Fine-tuning a pre-trained semantic segmentation model can be thought of as auto-tuning class-wise traversability values and adapting OOD representations caused by visual effects and terrain variations. Previously, these processes, such as class-wise traversability value tuning and semantic segmentation data labeling, were carried out manually[1, 15, 16].
The entire model parameters are fine-tuned via gradients derived from L2 loss for non-zero traversability labels and L1 loss with a weighting factor of 0.05 for zero traversability labels. Zero traversability labels are less penalized than non-zero traversability labels because, in our annotation pipeline, zero traversability labels may sometimes include both non-traversable and traversable but not yet traversed areas. The corresponding consideration in the loss function, as well as the model’s weak inductive bias from being initialized with the pre-trained semantic segmentation model, allow for robust traversability estimator learning in the presence of minor data noise. Random image augmentations (e.g., flipping, rotation, resize-cropping, and color jittering) are performed during training to deploy the semantic traversability estimator trained on egocentric videos to a variety of camera viewpoints.
III-D Usage for real-world robot navigation
[…]
IV Experimental Results
The proposed method for semantic traversability estimation was thoroughly evaluated, both qualitatively and quantitatively.
IV-A Evaluation of data collection and automatic annotation
We collected a wide range of egocentric videos and tested whether our automatic annotation method can identify areas that are well-aligned with humans’ visual preferences while also demonstrating robust performance across visually diverse video scenarios. The videos were taken in several countries (i.e., South Korea, Japan, and Italy) and cities (i.e., Daejeon, Seoul, Tokyo, Rome, and Venice), and included a variety of semantically traversable regions found in both urban and field environments, such as sidewalks with varying patterns, pedestrian areas, crosswalks, stairs, grassy terrains, snow regions, dirt areas, and others that are visually appealing to traverse but difficult to categorize. This diverse data collection could be done effectively due to the minimal hardware requirement of a single camera rather than the target robot itself with well-calibrated sensor settings. The efficacy of egocentric videos further highlights the potential for scalable dataset construction for semantic traversability estimation.
Figure 3: Automatic annotation results (KR-D: Daejeon, South Korea / KR-S: Seoul, South Korea / JP-T: Tokyo, Japan / IT-R: Rome, Italy / IT-V: Venice, Italy)
Figure 4: Stepwise annotation results. In Mask2former predictions, only sidewalks, crosswalks, roads, and buildings are color-segmented with pink, brown, purple, and blue if detected. In the final labels, sky-blue and red each indicates a value of 0.25 and 1.
The annotated results are shown in Figure 3. Clear traversable area labels consistent with humans’ semantic traversability were obtained by expanding the volunteer’s future footsteps using SAM. These expanded area labels, as opposed to point labels as in the previous work[2], provide rich supervision signals when training the semantic traversability estimator, and will be further explained in Section IV-B. Owing to the SAM’s highly generalizable segmentation capability, our automated annotation pipeline showed robust performance in a variety of video scenarios in both field and urban settings. The use of a large semantic segmentation model (i.e., Mask2former[14]) in conjunction with SAM resulted in improved final labels that complemented each other (Figure 4). SAM is particularly good at segmenting areas that include all footsteps based on visual cues, but it has limited performance in segmenting crosswalks where footsteps are placed not only on the white patterns but also on the road, resulting in overly large area labels. Mask2former, on the other hand, can reliably predict category-level semantic information, particularly for dominant categories such as ”road,” but not for non-dominant categories. Through the combination of these two models, we were able to obtain final annotations for images where either solely SAM cannot generate (e.g., crosswalk) or Mask2former cannot accurately predict (e.g., stairs, sidewalk, and pedestrian areas), and further achieve labels that include semantic category-level preference (i.e., road is less preferred than other traversed areas).
IV-B Evaluation of the trained semantic traversability estimator
We obtained the semantic traversability estimator by fine-tuning a lightweight semantic segmentation model (i.e., Segformer[13]) with approximately 2 hours and 20 minutes of egocentric videos. Videos were taken inside the KAIST campus (site size approximately 1.1M $\mathrm{m}^{2}$, 16.3 times the size of a soccer field), which is the testing region for our robot to perform autonomous navigation. The proposed automatic annotation pipeline produced 57K data tuples from the collected videos, each tuple containing an RGB image and pixel-wise traversability values, which were then randomly divided into train and validation sets with ratios of 0.9 and 0.1. The model was trained for 100 epochs with a batch size of 48 and a learning rate of 1e-5.
TABLE I: Quantitative evaluation results of semantic traversability estimators
Computation time [$\mathrm{m}\mathrm{s}$] # of parameters Same viewpoint Different viewpoint
Precision Recall IoU RMSE Precision Recall IoU RMSE
[Segformer + 1x1 conv] + Area labels (Ours) 14 / 61 3.7M
0.951
(0.927 / 0.975)
0.956
(0.972 / 0.940)
0.903 0.206
0.965
(0.953 / 0.977)
0.963
(0.975 / 0.950)
0.930 0.198
[Segformer + 1x1 conv] + Point labels [2] 14 / 61 3.7M
0.818
(0.985 / 0.650)
0.629
(0.260 / 0.997)
0.257 0.549
0.787
(0.989 / 0.584)
0.609
(0.220 / 0.997)
0.219 0.607
Autoencoder [22] 28 / 97 0.6M
0.610
(0.534 / 0.685)
0.620
(0.529 / 0.711)
0.358 0.596
0.622
(0.579 / 0.666)
0.628
(0.632 / 0.624)
0.437 0.598
Segformer + Heuristic [1, 15, 16] 94 / 886 3.7M
0.850
(0.750 / 0.950)
0.829
(0.918 / 0.741)
0.700 0.403
0.835
(0.752 / 0.918)
0.810
(0.889 / 0.730)
0.696 0.427
Mask2former + Heuristic [1, 15, 16] 89 / 692 43.9M
0.911
(0.861 / 0.961)
0.917
(0.963 / 0.871)
0.834 0.328
0.900
(0.843 / 0.957)
0.888
(0.945 / 0.832)
0.812 0.344
[Segformer (scratch) + 1x1 conv] + Area labels 14 / 61 3.7M
0.859
(0.745 / 0.974)
0.871
(0.975 / 0.767)
0.730 0.327
0.885
(0.793 / 0.977)
0.880
(0.982 / 0.777)
0.781 0.297
[Segformer (freeze) + 1x1 conv] + Area labels 14 / 61 3.7M
0.817
(0.656 / 0.977)
0.809
(0.985 / 0.633)
0.649 0.367
0.830
(0.687 / 0.972)
0.802
(0.989 / 0.615)
0.681 0.385
[Segformer (freeze) + U-Net] + Area labels 14 / 63 5.7M
0.918
(0.858 / 0.979)
0.933
(0.978 / 0.888)
0.843 0.239
0.942
(0.903 / 0.981)
0.942
(0.978 / 0.905)
0.885 0.207
[MobileSam] + Area labels 14 / 103 10.1M
0.927
(0.876 / 0.977)
0.934
(0.975 / 0.893)
0.858 0.230
0.947
(0.914 / 0.980)
0.941
(0.979 / 0.904)
0.897 0.200
  • The left and right values in ”Computation time” are each measured on a standard desktop GPU (NVIDIA GeForce RTX 4090) and an onboard embedded GPU (NVIDIA Jetson Orin). The left and right values inside the brackets in ”Precision” and ”Recall” represent traversable and non-traversable regions, respectively. The final ”Precision” and ”Recall” values are calculated by averaging the two values.
The trained semantic traversability estimator’s performance is compared to several prior methods demonstrated below:
  • [Segformer + 1x1 conv] + Point labels: The semantic segmentation model is fine-tuned with point labels obtained by projecting footsteps as in Wellhausen et al.[2]. Each footstep is converted to a circle with a radius of 30 pixels.
  • Autoencoder: Following the work by Schmid et al.[22], a denoising autoencoder modeled with convolution neural networks is trained to only reconstruct the labeled areas obtained from SAM. Areas with low reconstruction error (MSE loss smaller than 0.05) are then predicted as traversable.
  • Segformer + Heuristic: Lightweight semantic segmentation model is directly used with heuristic category-wise traversability values 111[Sidewalk, Path, Floor]: 1, [Grass, Sand, Hill, Dirt track, Land, Earth, Field]: 0.5, [Road]: 0.25 as in [1, 15, 16].
  • Mask2former + Heuristic: Large semantic segmentation model is directly used with heuristic category-wise traversability values 222[Pedestrian Area, Sidewalk, Lane Marking - Crosswalk]: 1, [Sand, Snow, Terrain]: 0.5, [Road, Lane Marking - General]: 0.25 as in [1, 15, 16].
We also compared the performance with variations of our method. Specifically, the model backbone was trained from scratch, the model backbone was frozen and only the adaptor was trained, the adaptor architecture was converted to a two-layer U-Net[28], and the model backbone was converted to MobileSAM[29], a compressed version of SAM through knowledge distillation. MobileSAM was chosen as an alternative model backbone over SAM because SAM requires more VRAM to fine-tune than the ones we have (NVIDIA GeForce RTX 4090 (24GB)) and is not suitable for real-time inference on an onboard embedded GPU due to its high computational cost.
[…]
Figure 5: Estimation results of semantic traversability estimators
Table I and Figure 5 present the quantitative and qualitative evaluation results, respectively. Our traversability estimator, which was obtained by fine-tuning a lightweight semantic segmentation model with area labels, outperformed other methods across all metrics. It demonstrated the highest accuracy in determining traversable areas, taking into account the highest precision, recall, and IoU values, as well as the lowest RMSE. Furthermore, it demonstrated the shortest computation time, allowing for semantic traversability estimation at a high frequency of 71Hz on a desktop GPU and 16Hz on an onboard embedded GPU. The model trained with only point labels, as in [2], showed low IoU values and only local area predictions due to small supervision signals during training, even with strong data augmentation. This demonstrates the significance of our automatic annotation method, which expands point labels to area labels through SAM, for more efficient training of image data with high visual diversity. Reconstruction-based methods [22] showed relatively low performance with large incorrect predictions due to the it’s limited capability to reconstruct high-frequency information (e.g., sidewalk patterns) and discriminate traversable regions based on reconstruction errors. Directly obtaining traversability values from semantic segmentation models showed lower accuracy due to incorrect predictions for non-dominant categories in the large-scale dataset the models were originally trained on, as well as high computation time due to a large number of neural network parameters and the computationally expensive argmax operation required for category-wise value assignments.
We also demonstrate the effectiveness of our training methodology, which involves leveraging pre-trained semantic segmentation model weights and fine-tuning them end-to-end. As shown in Table I, models that do not use pre-trained weights or only fine-tune the added layers showed lower accuracy regardless of the added layers’ size and neural network architecture. Interestingly, the model initialized with a segmentation model (i.e., MobileSAM [29]) showed lower accuracy than ours initialized with a semantic segmentation model (i.e., Segformer [13]). This implies that pre-trained representations of the semantic segmentation model are more suitable for semantic traversability estimation than those in the segmentation model and that fine-tuning them end-to-end is critical for adapting the OOD RGB features and terrain classes.
Figure 6: Estimation results for different camera viewpoints. All cameras are mounted closer to the ground than the camera mounted on a person’s chest. (A: GoPro camera, B: Single RGB-D camera configured horizontally, C: Two RGB-D cameras configured vertically)
The results in the ”Different viewpoint” section of Table I show that our semantic traversability estimator performed well and consistently on the test dataset containing different camera viewpoint images. This is also shown qualitatively in Figure 6-A. The trained semantic traversability estimator was further qualitatively evaluated using data from RGB-D cameras (Intel RealSense D435) with different image resolutions from the GoPro camera (Realsense: 640x360, GoPro: 1920x1080). Two camera configurations were tested: a horizontal configuration with a single camera facing forward and a vertical configuration with two cameras looking diagonally. As shown in both the image space predictions and projected 2.5D terrain maps in Figure 6, our model’s predictions for all these camera configurations are highly well-aligned with humans’ visual preferences. Overall, our experimental results show that our semantic traversability estimator can be generalized across different camera viewpoints, even when trained with egocentric videos, thanks to the strong data augmentation and weak inductive bias introduced from pre-trained model weights.
IV-C Evaluation of real-robot navigation
We tested our semantic traversability estimator in real-robot navigation scenarios, shown in Figure 7. The navigation experiments were conducted in several environments including sidewalks and narrow paths, where accurate traversability estimation and robust path planning are important. The quadrupedal robot Raibo2 found a path to navigate to the given goal destination considering both geometric and semantic traversability. Additionally, we performed obstacle avoidance in environments containing humans and untraversable boxes. The semantic traversability estimator identified them as untraversable objects and facilitated obstacle-overcoming path generation in conjunction with the geometric traversability.
Figure 7: Path planning with traversability estimation in real-world robot deployments. The green lines are the planned paths.
V Conclusion
We have proposed an effective methodology for training semantic traversability estimator with a dataset constructed with egocentric videos and an automated annotation strategy. We show that leveraging egocentric videos for robot navigation allows for scalable and efficient data acquisition. Extensive quantitative and qualitative experiments demonstrate that the yielded traversability estimator is highly accurate, camera viewpoint and configuration generalizable, and lightweight. Further deployment for autonomous quadruped robot navigation in an urban environment showcases the method’s capability for real-world applications.
Promising future directions include training a global semantic traversability estimator by scaling the dataset (e.g., worldwide data collection, leveraging internet-scale dataset[26]) and improving the data noise handling by incorporating representation-based methods[4, 30].